OpenStudy (anonymous):

Solve the system by graphing. y = 4/3x + 3 x2 + (y ̶ 3)2 = 25 A. (-3, 7) and (-1, 3) B. (-3, 7) and (3, -1) C. (7, -3) and (-1, 3) D. (7, -3) and (3, -1) I get (3,7) and (-3, -1)

OpenStudy (ybarrap):

Here are your points of intersection:  One way to check your answer is to see if when x=3 that y =7 for BOTH equations: checking 1st equation $$ y=\cfrac{4}{3}(3)+3=7\\ $$ checking 2nd equation $$ 3^2+(7-3)^2=9+16=25 $$ This checks out for both. Take the next point x=-3 and see if y=-1 $$ y=\cfrac{4}{3}(-1)+3=-1\\ $$ checking 2nd equation $$ 3^2+(-1-3)^2=9+16=25 $$ So both equations match. Your answer is therefore correct.